# SpecEdge: Scalable Edge-Assisted Serving Framework for Interactive LLMs

### v1

@@ Appendix A Broader Impact
Appendix B Limitations
SpecEdge is designed with the possibility that users bring their own GPUs for the edge drafting phase of their requests. Our cost-efficiency calculations assume that operators rent consumer-grade GPUs from edge cloud providers, but in principle, user-owned GPUs could further reduce costs. However, that scenario introduces additional complexities such as fault tolerance and security concerns if potentially malicious users operate their hardware. Although SpecEdge supports a distributed multi-user environment, addressing these broader challenges lies beyond the scope of this work.
Appendix C Additional Details on Evaluation Setup
For both SpecEdge and the server-only speculative decoding baseline, we adopt a tree construction algorithm from recent tree-based speculative decoding [Chen et al., 2024, Svirschevski et al., 2024]. Given a specified tree size, each forward pass of the draft model generates multiple parallel candidate tokens, which are then pruned based on cumulative log probabilities so that the total number of tokens remains within the tree budget. The total number of forward passes (the draft tree depth) is set as a hyperparameter. In our main experiments, we use the draft tree size to 32 for each request. For the server-only baseline, we find and select the optimal draft tree depth through exhaustive search, while SpecEdge determines its draft depth based on verification and network latencies as described in Section 4.3.
Throughout our evaluation, we use off-the-shelf target LLMs and draft models without any additional fine-tuning. For the main experiments, we employ Qwen3 32B/14B as the target LLM and Qwen3 1.7B/0.6B as the draft models, using the SpecBench dataset. To demonstrate broader applicability, the appendix includes experiments with Vicuna 33B and Llama2-13B-Chat-hf as target LLMs, alongside Sheared Llama 1.3B, Tiny Llama 1.1B, and JackFram 160M as draft models, evaluated on C4, OAsst, WikiText-2, and MTBench. For each query, we generate up to 256 output tokens. Because the prefill stage is identical for both the server-only baseline and SpecEdge, metrics such as time to first token remain the same; therefore, our reported metrics focus on the output tokens after prefill. Unless otherwise specified, SpecBench (spanning six different tasks) is used throughout the experiments.
@@ Appendix C Additional Details on Evaluation Setup
Vicuna 33B /SL 1.3B C4 3.47±$\pm 1.41$ 4.41±$\pm 2.03$ 28.12 49.96 (1.78x) 24.99 38.23 (1.53x)
OAsst 3.55±$\pm 1.37$ 4.35±$\pm 2.06$ 28.72 49.94 (1.74x) 25.53 37.89 (1.48x)
WikiText-2 3.42±$\pm 1.45$ 4.19±$\pm 1.98$ 27.26 47.94 (1.76x) 24.23 36.42 (1.50x)
MTBench 3.62±$\pm 1.72$ 4.57±$\pm 2.13$ 29.62 52.70 (1.78x) 26.33 40.21 (1.53x)
Llama2 13B /TL 1.1B C4 3.40±$\pm 1.35$ 3.48±$\pm 1.34$ 36.80 70.76 (1.92x) 32.72 53.72 (1.64x)
OAsst 3.57±$\pm 1.40$ 3.67±$\pm 1.39$ 38.55 69.49 (1.80x) 34.27 52.75 (1.54x)
WikiText-2 3.52±$\pm 1.37$ 3.61±$\pm 1.40$ 37.69 70.21 (1.86x) 33.50 53.30 (1.59x)
MTBench 3.67±$\pm 1.43$ 3.75±$\pm 1.46$ 40.96 76.69 (1.87x) 36.41 58.21 (1.60x)
Llama2 13B /JF 160M C4 3.03±$\pm 1.34$ 3.13±$\pm 1.37$ 42.72 66.20 (1.55x) 37.98 50.36 (1.33x)
OAsst 2.79±$\pm 1.37$ 2.85±$\pm 1.37$ 39.19 56.52 (1.44x) 34.84 42.95 (1.23x)
WikiText-2 2.72±$\pm 1.37$ 2.72±$\pm 1.39$ 37.87 55.39 (1.46x) 33.66 42.08 (1.25x)
MTBench 2.80±$\pm 1.39$ 2.80±$\pm 1.43$ 40.80 60.42 (1.48x) 36.42 45.92 (1.26x)
@@
Appendix D Performance Analysis with Diverse Models and Datasets
In Table 2, we showcase SpecEdge ’s adaptability across different target–draft model pairs and datasets, using a batch size of 1. We measure server throughput and cost efficiency gains, as well as the speedup achieved in inter-token latency. Table 3 compares these speedups for both SpecEdge and a server-only speculative decoding baseline against an autoregressive decoding reference.
Appendix E Performance with Non-Tree-based Speculative Decoding Method
To demonstrate the versatility of our approach beyond tree-structured methods, we implemented the speculative decoding technique from [Leviathan et al., 2023], which uses a linear candidate sequence rather than exploring multiple branching paths. This implementation allows us to evaluate whether SpecEdge’s core innovation—disaggregating drafting and verification between edge and server—generalizes effectively across different speculative decoding paradigms.
Tables 4 and 5 present the results of this evaluation, comparing SpecEdge against the conventional server-only deployment. The throughput measurements in Table 4 demonstrate SpecEdge’s efficiency advantages, while Table 5 quantifies the relative speedup over server-only autoregressive decoding. These results confirm that our disaggregated architecture delivers consistent improvements regardless of the underlying speculative decoding strategy, reinforcing the broad applicability of our approach.
@@ Appendix E Performance with Non-Tree-based Speculative Decoding Method
Appendix F Case Study: Proactive Edge Drafting in Action
@@
Experimental Setup. For this demonstration, we use a sample from the OAsst dataset with the query "What is a Dyson Sphere?". We employ Llama-3.2-3B [Grattafiori et al., 2024] as the target model and Llama-3.2-1B as the draft model.

### v2

@@ 6 Conclusion
Limitation and Future Work
SpecEdge is designed with flexibility in mind, supporting scenarios where users may contribute their own GPUs for the edge drafting phase. Our cost-efficiency analysis focuses on deployments using consumer-grade GPUs rented from edge cloud providers, but user-owned GPUs could further enhance cost savings and scalability. Extending SpecEdge to fully leverage user-operated hardware opens up exciting opportunities for decentralized and community-driven inference. At the same time, such scenarios raise new challenges in areas such as fault tolerance and security when untrusted devices participate in computation. While SpecEdge already supports a distributed multi-user environment, exploring these broader system and security aspects is an important direction for future work.
Acknowledgment
We thank the anonymous reviewers for providing helpful feedback and suggestions to improve our work. This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) of the Korea government (MSIT) (No. RS-2024-00398157).
Appendix A Additional Details on Evaluation Setup
For both SpecEdge and the server-only speculative decoding baseline, we adopt a tree construction algorithm from recent tree-based speculative decoding [chen2024sequoia, svirschevski2024specexec]. Given a specified tree size, each forward pass of the draft model generates multiple parallel candidate tokens, which are then pruned based on cumulative log probabilities so that the total number of tokens remains within the tree budget. The total number of forward passes (the draft tree depth) is set as a hyperparameter. In our main experiments, we use the draft tree size to 32 for each request. For the server-only baseline, we find and select the optimal draft tree depth through exhaustive search, while SpecEdge determines its draft depth based on verification and network latencies as described in Section 4.3.
Throughout our evaluation, we use off-the-shelf target LLMs and draft models without any additional fine-tuning. For the main experiments, we employ Qwen3 32B/14B as the target LLM and Qwen3 1.7B/0.6B as the draft models, using the SpecBench dataset. To demonstrate broader applicability, the appendix includes experiments with Vicuna 33B and Llama2-13B-Chat-hf as target LLMs, alongside Sheared Llama 1.3B, Tiny Llama 1.1B, and JackFram 160M as draft models, evaluated on C4, OAsst, WikiText-2, and MTBench. For each query, we generate up to 256 output tokens. Since the prefill stage in SpecEdge can involve parallel processing on both server and edge, the prefill latency is determined by $\max(\text{server\_time},\text{edge\_time})$ which makes metrics such as time to first token comparable to the baseline; therefore, our reported metrics focus on the output tokens after prefill. Unless otherwise specified, SpecBench (spanning six different tasks) is used throughout the experiments.
@@ Appendix A Additional Details on Evaluation Setup
[…]
WikiText-2 $\pm 1.37 2.72$$\pm 1.39$ 37.87 55.39 (1.46x) 33.66 42.08 (1.25x)
MTBench $\pm 1.39 2.80$$\pm 1.43$ 40.80 60.42 (1.48x) 36.42 45.92 (1.26x)
Appendix B Comprehensive Performance Analysis
B.1 Performance with Diverse Models and Datasets
In Table 2, we showcase SpecEdge ’s adaptability across different target–draft model pairs and datasets, using a batch size of 1. We measure server throughput and cost efficiency gains, as well as the speedup achieved in inter token latency. Table 3 compares these speedups for both SpecEdge and a server-only speculative decoding baseline against an autoregressive decoding reference.
@@ B.1 Performance with Diverse Models and Datasets
B.2 Performance with Lighter GPUs
We use the RTX 4090 as representative of consumer-grade GPUs, now widespread in both edge-cloud providers and user devices. However, to demonstrate broader generalizability, we conducted additional experiments with lighter consumer-grade GPUs. We measured SpecEdge performance using the RTX 3060 Ti and the RTX 2080 Ti. Table 4 shows the inter token latency and throughput with target/draft model pairs configured as Qwen3-14B/1.7B and Qwen3-14B/0.6B. Compared to the server-only approach, SpecEdge still attains meaningful throughput improvements even with less powerful GPUs.
Table 4: Performance comparison of lighter edge GPUs.
Target/Draft Model Edge GPU Peak FP16 TFLOPS Memory Bandwidth (GB/s) Inter token Latency (ms) Server Throughput (tok/s)
Qwen3-14B/1.7B RTX 3060 Ti 16.20 448.0 36.818 50.297
RTX 2080 Ti 26.90 616.0 34.409 54.657
Server-only (A100 40GB) 312 1555 32.451 30.816
Qwen3-14B/0.6B RTX 3060 Ti 16.20 448.0 36.818 56.135
RTX 2080 Ti 26.90 616.0 34.409 54.657
Server-only (A100 40GB) 312 1555 32.326 30.935
B.3 Performance with Non-Tree-based Speculative Decoding Method
To demonstrate the versatility of our approach beyond tree-structured methods, we implemented the speculative decoding technique from [leviathan2023fast], which uses a linear candidate sequence rather than exploring multiple branching paths. This implementation allows us to evaluate whether SpecEdge’s core innovation—disaggregating drafting and verification between edge and server—generalizes effectively across different speculative decoding paradigms.
Tables 5 and 6 present the results of this evaluation, comparing SpecEdge against the conventional server-only deployment. The throughput measurements in Table 5 demonstrate SpecEdge’s efficiency advantages, while Table 6 quantifies the relative speedup over server-only autoregressive decoding. These results confirm that our disaggregated architecture delivers consistent improvements regardless of the underlying speculative decoding strategy, reinforcing the broad applicability of our approach.
@@ B.3 Performance with Non-Tree-based Speculative Decoding Method
B.4 Performance with Batch Drafting Method
While our main deployment architecture assumes that each concurrent request utilizes a dedicated edge GPU, we investigate an alternative architecture where a single edge GPU generates draft tokens for multiple concurrent requests. This alternative method offers trade-off for scenarios where operators want to utilize fewer consumer-grade edge GPUs. We evaluate this configuration by batching multiple requests on the RTX 4090 drafter with batch sizes from 2 to 4, measuring both inter token latency and cost efficiency across different target/draft model pairs.
The results show a trade-off between cost efficiency and latency. As shown in Table 7, cost efficiency improves with the alternative method, ranging from 4.4% to 29.5% better across different configurations due to better edge GPU utilization. However, this comes at the expense of increased inter token latency (5.9% to 19.0% slower), primarily caused by contention from batching multiple requests and longer draft-to-verify cycles. This alternative deployment method could be preferable in cost-sensitive scenarios where higher latency is acceptable.
Table 7: Performance comparison of alternative deployment method.
Target/Draft Model Batch Size Inter Token Latency (ms) Cost Efficiency (1k toks/\$)
Qwen3-14B/1.7B 2 31.777 (6.8% slower) 67.299 (14.7% better)
3 36.617 (13.0% slower) 71.369 (29.5% better)
4 40.099 (16.7% slower) 85.662 (24.5% better)
Qwen3-14B/0.6B 2 30.494 (8.5% slower) 67.680 (5.8% better)
3 34.028 (15.2% slower) 80.252 (12.0% better)
4 39.035 (19.0% slower) 82.376 (20.0% better)
Qwen3-32B/1.7B 2 43.539 (5.9% slower) 51.245 (4.4% better)
3 48.889 (3.9% slower) 59.635 (21.3% better)
4 58.402 (9.5% slower) 57.857 (25.1% better)
B.5 Performance under Reasoning Mode
Many modern LLMs support reasoning capabilities to enhance their inference performance. Reasoning tokens often exhibit highly repetitive patterns, which can impact the acceptance rate in speculative decoding. To investigate the inference performance with reasoning capabilities, we measured the accepted tokens, server throughput, and cost efficiency with and without reasoning mode enabled.
Table 8 presents the accepted tokens, server throughput, and cost efficiency for each target-draft model pair. When generating reasoning tokens, we observe consistent improvements across all three metrics: accepted tokens, server throughput, and cost efficiency. These results indicate that SpecEdge can inherently benefit from improved speculative decoding performance coming from redundancy in reasoning processes.
Table 8: Performance comparison of SpecEdge with and without reasoning mode.
Accepted Tokens Server Throughput (tok/s) Cost Efficiency (1k toks/\$)
Target/Draft Model Non-Reasoning Reasoning Non-Reasoning Reasoning Non-Reasoning Reasoning
Qwen3-14B/1.7B $3.80\pm 1.54$ $4.17\pm 1.39$ 64.343 72.249 48.883 54.898
Qwen3-14B/0.6B $3.68\pm 1.50$ $4.10\pm 1.36$ 70.703 81.890 53.693 62.157
Qwen3-32B/1.7B $4.09\pm 1.95$ $4.62\pm 1.90$ 24.880 27.617 39.430 43.373
Appendix C Cost Efficiency with Various GPU Providers and Cloud Service Providers
We anchored our cost estimates to widely available public pricing: A100 GPUs from Google Cloud Platform and RTX 4090 GPUs from Vast.ai [Vastai]. To validate the robustness of our cost-efficiency claims, we conducted additional experiments across multiple cloud environments, including various GPU providers (Vast.ai, Runpod [runpod], and TensorDock [TensorDock]) and major cloud service providers (Google Cloud Platform [GCP], Amazon Web Services [AWS], and Microsoft Azure [Azure]). Table 10 and 10 show the GPU pricing from different providers. Table 11 demonstrates that SpecEdge maintains cost efficiency improvements consistently across all tested configurations.
Table 9: Edge GPU Pricing.
GPU GPU Provider Cost (\$/hr)
RTX 4090 RunPod 0.69
Vast.ai 0.35
TensorDock 0.359
RTX Pro 6000 Vast.ai 1.08
TensorDock 1.15
Table 10: Server GPU Pricing.
GPU Cloud Service Provider Cost (\$/hr)
A100 40GB Google Cloud Platform 4.05
Amazon Web Services 4.10
A100 80GB Google Cloud Platform 5.05
Amazon Web Services 5.12
Microsoft Azure 3.673
Table 11: Cost efficiency comparison of GPU Providers and Cloud Service Providers.
Cost Efficiency (1k toks/\$)
Target/Draft Model Server/Edge GPU GPU Provider Google Cloud Platform Amazon Web Service Microsoft Azure
Qwen3-14B/1.7B A100 40GB / RTX 4090 Vast.ai 51.018 ($1.87\times$) 50.485 ($1.87\times$) -
RunPod 44.721 ($1.64\times$) 44.311 ($1.65\times$) -
TensorDock 50.829 ($1.87\times$) 50.30($1.87\times$) -
Baseline (Server-only) 27.222 26.890 -
Qwen3-14B/1.7B A100 40GB / RTX Pro 6000 Vast.ai 47.68 ($1.75\times$) 47.30 ($1.76\times$) -
TensorDock 46.64 ($1.71\times$) 46.27 ($1.72\times$) -
Baseline (Server-only) 27.222 26.890 -
Qwen3-14B/0.6B A100 40GB / RTX 4090 Vast.ai 52.938 ($1.86\times$) 52.386 ($1.88\times$) -
RunPod 46.382 ($1.63\times$) 45.957 ($1.65\times$) -
TensorDock 52.741 ($1.86\times$) 52.192 ($1.88\times$) -
Baseline (Server-only) 28.415 27.802 -
Qwen3-32B/1.7B A100 80GB / RTX 4090 Vast.ai 31.619 ($1.82\times$) 31.332 ($1.83\times$) 41.757 ($1.75\times$)
RunPod 28.453 ($1.64\times$) 28.144 ($1.65\times$) 36.280 ($1.52\times$)
TensorDock 31.619 ($1.82\times$) 31.239 ($1.83\times$) 41.591 ($1.75\times$)
Baseline (Server-only) 17.330 17.090 23.822
Appendix D Case Study: Proactive Edge Drafting in Action
@@
Experimental Setup. For this demonstration, we use a sample from the OAsst dataset with the query "What is a Dyson Sphere?". We employ Llama-3.2-3B [grattafiori2024llama] as the target model and Llama-3.2-1B as the draft model.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction include the paper’s main claims and contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are discussed in the appendix, limitations section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include Theorems or Lemmas that need formal proofs. The formulas are well-numbered and clearly state their notations.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The end-to-end system and methods are thoroughly described. The models, dataset, and hyperparameters used for the evaluation are provided in detail.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We submit our code as supplementary material, which is fully reproducible. We will publicly open the submitted code after the review.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All the experiment settings and methods are specified in detail.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The error bars are provided in the evaluation figures and tables.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We specify the type of compute resources used for each experiment setting.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The potential societal impacts are discussed in the appendix, broader impacts section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper suggests a new serving system paradigm that is orthogonal to the release of data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The models, data, and code we used or reproduced are properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code for the end-to-end system in the paper is submitted as an anonymized zip file through the supplementary material. The code is provided with documentation with instructions to run the code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.